# Data Release 1 of the Dark Energy Spectroscopic Instrument

### v1

@@ 3.2 Sample Distributions
In Figure 6 we show the redshift distribution of stellar (left-hand panel) and extragalactic (right-hand panels) targets in DR1 with well-measured redshifts. In each panel, the gray histogram shows the distribution of objects spectroscopically classified by Redrock (RR) into stars, galaxies, and quasars (SPECTYPE=STAR, GALAXY, and QSO; see §3.1.2). The light-blue, dark-green, red, dark-blue, and purple histograms show the redshift distribution of MWS, BGS, LRG, ELG, and QSO tracers, respectively. As discussed in DESI Collaboration et al. (2024a, b) and the collection of papers listed in Table 3, these redshift distributions match the finely-tuned target-selection algorithms developed for each class of DESI target.
@@
Table 7 summarizes the exact number of unique, well-measured redshifts in DR1. In this table we show the detailed breakdown of the number of good redshifts as a function of target class (including secondary, SCND, targets; see §§1.1 and 2.2), and the full set of observational programs included in DR1 (see §2.3). In addition, the last column gives the total number of good redshifts within each program, although we emphasize that this total cannot be derived from the sum of the preceding columns because it includes objects which may belong to one or more (or none) of the listed target classes.
Finally, in Figure 7 we show examples of coadded spectra for each of the primary extragalactic target classes.
3.3 Data Products
@@
In this section we provide a high-level overview of the organizational (directory) structure of the data included in DR1 (see §4 for information regarding how the data can be accessed). For a complete description of all the directories and files in DR1, as well as their provenance and inter-dependencies, please see the DESI data model.171717https://desidatamodel.readthedocs.io
@@
Table 8 shows key elements of the DR1 file and directory structure and relative locations, and the number of files and total size in terabytes. Most users will be interested in the data and catalogs in spectro/redux/iron/181818As discussed in §1.2, Iron is the primary spectroscopic production for DR1., the LSS catalogs in survey/catalogs/dr1/LSS/ (see §3.4), and the VACs in vac/dr1/ (see §3.5 and Appendix F).
@@
In the top-level production directory, spectro/redux/iron/tiles-iron.fits contains a catalog of all DESI tiles included in Iron. This file can be used to quickly assess the observational footprint of DR1, and to filter the set of tiles by survey (SURVEY) or program (PROGRAM). Because tiles may be observed on multiple exposures spanning multiple nights (see §2.3), more detailed per-exposure information can be found in the spectro/redux/iron/exposures-iron.fits file, for example, for time-domain studies or for comparisons of systematic differences of data obtained on different nights.
@@
In §3.3.1 we describe the organization of the spectra, coadds, and lower-level (per-observation) redshift files, and in §3.3.2 we describe the merged redshift catalogs, which are joined across different combinations of DESI surveys and programs. In addition, §3.3.3 describes specific updates to the parent target catalogs used for DR1 observations. For additional information and descriptions of the other kinds of files available in DR1, such as the fiber assignment and parent photometric-target catalogs, please see DESI Collaboration et al. (2024b) and the DESI data model.
@@
3.3.1 Spectra, Coadds, and Redshifts
Per-exposure spectra, coadded spectra (coadds), redshifts, and after-burner catalogs are organized into two broad groups or categories: per-tile and “healpix”. Tile-based spectra can be found in spectro/redux/iron/tiles/; these spectra combine information across multiple exposures of the same tile, but not across different tiles (even if the same target was observed on multiple tiles; see §1.2). Meanwhile, healpix coadds combine all the available exposures of targets observed on different tiles into a given HEALPix pixel on the sky (nested scheme, using NSide=64; Górski et al., 2005; Zonca et al., 2019); healpix coadds can be found under spectro/redux/iron/healpix/.
In the case of both the tile-based and healpix coadds, data are not combined across surveys (SV1, SV2, SV3, main, special) or programs (bright, dark, backup). This decision is driven by the desire to prioritize the uniformity of the data in each survey/program combination (but see the ZCAT_PRIMARY flag described in §3.3.2). In general, we expect most users to utilize the HEALPix-grouped spectra and redshifts, while the tile-based results will be used for more specialized (including cosmological) analyses.
We organize healpix-grouped outputs in subdirectories by HEALPix number; however, in order to avoid having tens of thousands of subdirectories at the same level, these subdirectories are additionally grouped by integers given by int(healpix/100). For example, all SURVEY=main, PROGRAM=dark data for targets in nested NSide=64 HEALPix number 31542 can be found in spectro/redux/iron/healpix/main/dark/315/31542/. The files in this directory include the per-exposure (uncoadded) spectra, coadded spectra (across exposures but not across cameras), Redrock redshift fits and classifications, and the outputs from the pipeline after-burner algorithms (described in §3.1.2).
@@ 3.3.1 Spectra, Coadds, and Redshifts
3.3.2 Merged Redshift Catalogs
Merged redshift catalogs, combined across thousands of smaller, individual files and catalogs, can be found in spectro/redux/iron/zcatalog/v1/. Like the spectra and coadds described in §3.3.1, these catalogs come in multiple groups, for example, combining all the cumulative, tile-based redshifts for a given survey and program into a single file, with different files for different survey/program combinations. Future data releases may include v2 (or higher) version numbers with a modified data model optimized for the increasingly large catalogs. In general, we recommend using the highest version number available, which is v1 in the case of DR1.
@@ 3.3.2 Merged Redshift Catalogs
3.3.3 Parent Target Catalogs
Target catalogs used as input for DESI observations were previously published in the Early Target Selection (ETS) release described in Myers et al. (2023a) and included as part of the EDR.212121https://data.desi.lbl.gov/public/edr/target The photometric catalogs from which primary DESI targets have been selected remain unchanged relative to the EDR and can be found in the target/catalogs directory tree (see Table 8). Meanwhile, some catalogs of secondary222222https://data.desi.lbl.gov/public/dr1/target/catalogs/dr9/1.3.0 and calibration232323https://data.desi.lbl.gov/public/dr1/target/catalogs/dr9/2.2.0 targets have been updated between the ETS and DR1 in a manner consistent with the schemas described in §3.2, §4.3 and §4.4 of Myers et al. (2023a). Additional details regarding secondary targets can be found in Appendix C; in addition, some random catalogs (see §4.5.1 of Myers et al., 2023a) were repackaged into HEALPix (Górski et al., 2005) pixels to simplify LSS analyses, and these are also included as part of DR1.242424https://data.desi.lbl.gov/public/dr1/target/catalogs/dr9/2.4.0
3.4 Large-Scale Structure Catalogs
Together with the raw and intermediate DESI data, DR1 includes all the LSS catalogs needed to reproduce the DR1 clustering and cosmology analyses, including the BAO measurements (DESI Collaboration et al., 2024d), the full-shape analyses (DESI Collaboration et al., 2024f), and the derivation of the baseline cosmological constraints from the four extragalactic DESI tracers (BGS, LRG, ELG, and QSO; DESI Collaboration et al., 2025b). For complete details regarding the sample definitions and characteristics of the LSS clustering catalogs in DR1, see DESI Collaboration et al. (2024c) and Ross et al. (2025). Here, we briefly summarize the broad organizational structure of the LSS catalogs, and how they can be accessed (see also Table 8).
@@ 3.4 Large-Scale Structure Catalogs
| v1.2 | Baseline used in DR1 BAO analysis and cosmological results. | DESI Collaboration et al. (2024d, 2025b) | v1.2-DR1 |
| v1.5 | Minor bug fixes; should be used for all clustering measurements at scales greater than the fiber patrol radius (.) | DESI Collaboration et al. (2024f); Chaussidon et al. (2024) | v1.5-DR1 |
Unlike in the EDR, where the LSS catalogs were released as a VAC (see §4 of DESI Collaboration et al. 2024b), the LSS catalogs in DR1 can be found under the survey/catalogs/dr1/LSS directory (see §3.3 and Table 8). This directory contains all the auxiliary files defining the DR1 tiles, potential assignments for 18 random catalogs, and the input redshifts from the Iron spectroscopic production.
The ready-to-use clustering catalogs and associated randoms are available in survey/catalogs/dr1/LSS/iron/LSScats. In this directory, there are two distinct versions of these catalogs, v1.2 and v1.5 (see Table 9). A discussion of the differences between these versions can be found in Appendix B of DESI Collaboration et al. (2024c).
@@
Data containing the information on 128 alternative realizations of the DR1 fiber-assignment history (used, e.g., to determine pairwise inverse probability, PIP, weights for 2-point clustering measurements; see Lasker et al. 2025 and Bianchi et al. 2024), together with the catalog version v1.5pip described in Appendix C of DESI Collaboration et al. (2024c), will be published in the near future.
@@
3.5 Value-Added Catalogs
@@
| LS/DR9 Photometry | Merged targeting catalogs and Legacy Surveys DR9 Tractor photometric |
| catalogs for all observed and potential DESI targets. | |
| Sky Spectra | Exemplar sky spectra with detailed metadata from the DESI pipeline. |
@@
| MWS | Analysis of stellar spectra by the MWS Working Group. |
| MWS BHB | Catalog of spectroscopically confirmed blue horizontal branch (BHB) stars |
| (Byström et al., 2024). | |
| MWS SpecDis | Spectrophotometric distances for million stars in DR1 predicted using a neural network trained on stellar spectra (Li et al., 2025). |
@@
| Stellar Reddening | Spectra and catalog of stars used in dust reddening measurements (Zhou et al., 2024). |
@@
| DESI HETDEX | HETDEX and DESI spectra for Hobby-Eberly Telescope Dark Energy Experiment (HETDEX) Ly emitter candidates observed by DESI (Landriau et al., 2025). |
| DESIVAST | Cosmic voids identified within the DESI DR1 volume (Rincon et al., 2024). |
@@
| EmFit | Emission-line fitting results for galaxies (Pucha et al., 2024, Pucha et al. 2025, in preparation). |
| FastSpecFit | Spectrophotometric fitting results from the FastSpecFit stellar continuum and |
| emission-line modeling code. | |
| Extended Halo-based Group | Halo-based group catalog based on Legacy Surveys DR9 for -band apparent magnitude galaxies. |
| Mass EMLines | Stellar mass and emission line measurements for galaxies in DR1 (Zou et al., 2024). |
@@
| AGN/QSO Summary | AGN and QSO identifications for all target classes in DESI DR1 (Juneau et al. 2025, in preparation). |
| AGN Host Properties | Stellar masses and other physical properties from spectral energy distribution |
| modeling which includes AGN templates (Siudek et al., 2024). | |
| BHMass | Iron-corrected supermassive black hole masses based on Mg II at |
| (Pan et al., 2025). | |
| DLA NN and GP Finder | DLA parameters and detections using the NN and GP DLA finders (Wang et al., 2022; Ho et al., 2020). |
| DLA Template Finder | DLA parameters and detections using DLA Toolkit (Brodzeller et al. 2025, in preparation). |
| MgII Absorber | Summarized information of Mg II absorption systems in DESI quasars (Napolitano et al., 2023). |
| ZLyA | Updated redshifts and BAL information used in the Ly Y1 BAO analysis (DESI Collaboration et al., 2025a). |
@@
| Ly Forest Y1 Deltas | Measured flux-transmission field used in the Ly Y1 BAO analysis (DESI Collaboration et al., 2025a). |
| Ly Forest Y1 Correlations | Measured correlations, distortion matrices and covariances used for the |
| Ly Y1 BAO analysis (DESI Collaboration et al., 2025a). | |
Each DESI data release includes multiple VACs, which are additional data products, catalogs, and documentation contributed by members of the DESI science collaboration. These VACs are built upon the core data products (spectra, classifications, redshifts) from this data release, and include additional data useful for a variety of scientific analyses. DR1 includes a total of 24 distinct VACs at the date of the public data release, although additional VACs may be added to DR1 following this date. The most up-to-date list of VACs and their associated documentation and references can be found online at the public DESI website.272727https://data.desi.lbl.gov/doc/vac
4 Data Access
In this section we provide a brief overview of how to access DR1. At the time of this writing, DESI catalogs and spectra can be accessed via file download (§4.1) or a searchable database (§4.2). We demonstrate these access methods, as well as techniques for exploring and manipulating DESI data, through a variety of tutorials and examples (§4.3). The DESI collaboration may deploy additional data-access methods in the future, such as web services designed for interactive data exploration or interfaces enabling users to download individual spectra or custom collections of spectra. The latest information on how to access DESI data can always be found in the DESI documentation.282828https://data.desi.lbl.gov/doc/access

### v2

@@ III.2 Sample Distributions
In Figure 6 we show the radial velocity distribution of stellar targets (left-hand panel) and the redshift distribution of extragalactic targets (right-hand panels) in DR1 with well-measured redshifts. In each panel, the gray histogram shows the distribution of objects spectroscopically classified by Redrock (RR) into stars, galaxies, and quasars (SPECTYPE=STAR, GALAXY, and QSO; see §III.1.2). The light-blue histrogram shows the radial velocity distribution of MWS targets, and the dark-green, red, dark-blue, and purple histograms show the redshift distribution of BGS, LRG, ELG, and QSO tracers, respectively. As discussed in DESI Collaboration et al. (2024a, b) and the collection of papers listed in Table 3, these redshift distributions match the finely-tuned target-selection algorithms developed for each class of DESI target.
@@
Table 7 summarizes the exact number of unique, well-measured redshifts in DR1. In this table we show the detailed breakdown of the number of good redshifts as a function of target class (including secondary, SCND, targets; see §§I.1 and II.2), and the full set of observational programs included in DR1 (see §II.3). In addition, the last column gives the total number of good redshifts within each program, although we emphasize that this total cannot be derived from the sum of the preceding columns because it includes objects which may belong to one or more (or none) of the listed target classes.
Finally, in Figure 7 we show examples of coadded spectra and Legacy Surveys color images for each of the primary extragalactic target classes. BGS galaxies tend to include both obvious stellar continuum and emission lines in their spectra with a spatially resolved morphology. LRGs tend to have spectra dominated by a red stellar continuum indicative of old stellar populations and images characterized by red colors with a compact (e.g., de Vaucouleurs) morphology. ELGs have, on-average a fainter and bluer continuum with respect to the other galaxy classes, with distinct [O II] doublet emission and their images often appear faint, blue and barely resolved. Finally, QSO spectra are characterized by a blue continuum and broad emission lines arising from the nuclear accretion disk and surrounding ionized gas, and they all have a point-source morphology by selection.
III.3 Data Products
@@
In this section we provide a high-level overview of the organizational (directory) structure of the data included in DR1 (see §IV for information regarding how the data can be accessed). For a complete description of all the directories and files in DR1, as well as their provenance and inter-dependencies, please see the DESI data model.171717https://desidatamodel.readthedocs.io
@@
Table 8 shows key elements of the DR1 file and directory structure and relative locations, and the number of files and total size in terabytes. Most users will be interested in the data and catalogs in spectro/redux/iron/181818As discussed in §I.2, Iron is the primary spectroscopic production for DR1., the LSS catalogs in survey/catalogs/dr1/LSS/ (see §III.4), and the VACs in vac/dr1/ (see §III.5 and Appendix F).
@@
In the top-level production directory, spectro/redux/iron/tiles-iron.fits contains a catalog of all DESI tiles included in Iron. This file can be used to quickly assess the observational footprint of DR1, and to filter the set of tiles by survey (SURVEY) or program (PROGRAM). Because tiles may be observed on multiple exposures spanning multiple nights (see §II.3), more detailed per-exposure information can be found in the spectro/redux/iron/exposures-iron.fits file, for example, for time-domain studies or for comparisons of systematic differences of data obtained on different nights.
@@
In §III.3.1 we describe the organization of the spectra, coadds, and lower-level (per-observation) redshift files, and in §III.3.2 we describe the merged redshift catalogs, which are joined across different combinations of DESI surveys and programs. In addition, §III.3.3 describes specific updates to the parent target catalogs used for DR1 observations. For additional information and descriptions of the other kinds of files available in DR1, such as the fiber assignment and parent photometric-target catalogs, please see DESI Collaboration et al. (2024b) and the DESI data model.
@@
III.3.1 Spectra, Coadds, and Redshifts
Per-exposure spectra, coadded spectra (coadds), redshifts, and after-burner catalogs are organized into two broad groups or categories: per-tile and “healpix”. Tile-based spectra can be found in spectro/redux/iron/tiles/; these spectra combine information across multiple exposures of the same tile, but not across different tiles (even if the same target was observed on multiple tiles; see §I.2). Meanwhile, healpix coadds combine all the available exposures of targets observed on different tiles into a given HEALPix pixel on the sky (nested scheme, using NSide=64; gorski05a; zonca19a); healpix coadds can be found under spectro/redux/iron/healpix/.
In the case of both the tile-based and healpix coadds, data are not combined across surveys (SV1, SV2, SV3, main, special) or programs (bright, dark, backup). This decision is driven by the desire to prioritize the uniformity of the data in each survey/program combination (but see the ZCAT_PRIMARY flag described in §III.3.2). In general, we expect most users to utilize the HEALPix-grouped spectra and redshifts, while the tile-based results will be used for more specialized (including cosmological) analyses.
We organize healpix-grouped outputs in subdirectories by HEALPix number; however, in order to avoid having tens of thousands of subdirectories at the same level, these subdirectories are additionally grouped by integers given by int(healpix/100). For example, all SURVEY=main, PROGRAM=dark data for targets in nested NSide=64 HEALPix number 31542 can be found in spectro/redux/iron/healpix/main/dark/315/31542/. The files in this directory include the per-exposure (uncoadded) spectra, coadded spectra (across exposures but not across cameras), Redrock redshift fits and classifications, and the outputs from the pipeline after-burner algorithms (described in §III.1.2).
@@ III.3.1 Spectra, Coadds, and Redshifts
III.3.2 Merged Redshift Catalogs
Merged redshift catalogs, combined across thousands of smaller, individual files and catalogs, can be found in spectro/redux/iron/zcatalog/v1/. Like the spectra and coadds described in §III.3.1, these catalogs come in multiple groups, for example, combining all the cumulative, tile-based redshifts for a given survey and program into a single file, with different files for different survey/program combinations. Future data releases may include v2 (or higher) version numbers with a modified data model optimized for the increasingly large catalogs. In general, we recommend using the highest version number available, which is v1 in the case of DR1.
@@ III.3.2 Merged Redshift Catalogs
III.3.3 Parent Target Catalogs
Target catalogs used as input for DESI observations were previously published in the Early Target Selection (ETS) release described in Myers et al. (2023) and included as part of the EDR.212121https://data.desi.lbl.gov/public/edr/target The photometric catalogs from which primary DESI targets have been selected remain unchanged relative to the EDR and can be found in the target/catalogs directory tree (see Table 8). Meanwhile, some catalogs of secondary222222https://data.desi.lbl.gov/public/dr1/target/catalogs/dr9/1.3.0 and calibration232323https://data.desi.lbl.gov/public/dr1/target/catalogs/dr9/2.2.0 targets have been updated between the ETS and DR1 in a manner consistent with the schemas described in §3.2, §4.3 and §4.4 of Myers et al. (2023). Additional details regarding secondary targets can be found in Appendix C; in addition, some random catalogs (see §4.5.1 of Myers et al., 2023) were repackaged into HEALPix (gorski05a) pixels to simplify LSS analyses, and these are also included as part of DR1.242424https://data.desi.lbl.gov/public/dr1/target/catalogs/dr9/2.4.0
III.4 Large-Scale Structure Catalogs
Together with the raw and intermediate DESI data, DR1 includes all the LSS catalogs needed to reproduce the DR1 clustering and cosmology analyses, including the BAO measurements (DESI Collaboration et al., 2024f), the full-shape analyses (DESI Collaboration et al., 2024c), and the derivation of the baseline cosmological constraints from the four extragalactic DESI tracers (BGS, LRG, ELG, and QSO; DESI Collaboration et al., 2025a). For complete details regarding the sample definitions and characteristics of the LSS clustering catalogs in DR1, see DESI Collaboration et al. (2024d) and Ross et al. (2025). Here, we briefly summarize the broad organizational structure of the LSS catalogs, and how they can be accessed (see also Table 8).
@@ III.4 Large-Scale Structure Catalogs
| v1.2 | Baseline used in DR1 BAO analysis and cosmological results. | DESI Collaboration et al. (2024f, 2025a) | v1.2-DR1 |
| v1.5 | Minor bug fixes; should be used for all clustering measurements at scales greater than the fiber patrol radius (.) | DESI Collaboration et al. (2024c); Chaussidon et al. (2024) | v1.5-DR1 |
Unlike in the EDR, where the LSS catalogs were released as a VAC (see §4 of DESI Collaboration et al. 2024b), the LSS catalogs in DR1 can be found under the survey/catalogs/dr1/LSS directory (see §III.3 and Table 8). This directory contains all the auxiliary files defining the DR1 tiles, potential assignments for 18 random catalogs, and the input redshifts from the Iron spectroscopic production.
The ready-to-use clustering catalogs and associated randoms are available in survey/catalogs/dr1/LSS/iron/LSScats. In this directory, there are two distinct versions of these catalogs, v1.2 and v1.5 (see Table 9). A discussion of the differences between these versions can be found in Appendix B of DESI Collaboration et al. (2024d).
@@
Data containing the information on 128 alternative realizations of the DR1 fiber-assignment history (used, e.g., to determine pairwise inverse probability, PIP, weights for 2-point clustering measurements; see Lasker et al. 2025 and Bianchi et al. 2024), together with the catalog version v1.5pip described in Appendix C of DESI Collaboration et al. (2024d), will be published in the near future.
@@
III.5 Value-Added Catalogs
Each DESI data release includes multiple VACs, which are additional data products, catalogs, and documentation contributed by members of the DESI science collaboration. These VACs are built upon the core data products (spectra, classifications, redshifts) from this data release, and include additional data useful for a variety of scientific analyses. DR1 includes a total of 27 distinct VACs at the date of the public data release, although additional VACs may be added to DR1 following this date. The most up-to-date list of VACs and their associated documentation and references can be found online at the public DESI website.272727https://data.desi.lbl.gov/doc/vac
Table 10 provides a brief overview of the VACs released concurrently with DR1, along with short descriptions of the contents (and associated reference, if available) of each VAC. Additional details regarding each VAC can be found in Appendix F.
@@ III.5 Value-Added Catalogs
| LS/DR9 Photometry | Merged targeting catalogs and Legacy Surveys DR9 Tractor photometric catalogs for all observed and potential DESI targets. |
| Sky Spectra | Example sky spectra with detailed metadata from the DESI pipeline. |
| BAO Cosmology Results | Cosmology chains and posterior maximization results for the DESI DR1 BAO cosmology results (DESI Collaboration et al., 2025a). |
| Full Shape Cosmology Results | Cosmology chains and posterior maximization results for the DESI DR1 full shape analysis results (DESI Collaboration et al., 2024c). |
@@
| MWS | Analysis of stellar spectra by the MWS Working Group (koposov25a). |
| MWS BHB | Catalog of spectroscopically confirmed blue horizontal branch (BHB) stars (bystrom25a). |
| MWS SpecDis | Spectrophotometric distances for million stars in DR1 predicted using a neural network trained on stellar spectra (li25a). |
@@
| Stellar Reddening | Spectra and catalog of stars used in dust reddening measurements (zhou24a). |
@@
| DESI HETDEX | HETDEX and DESI spectra for Hobby-Eberly Telescope Dark Energy Experiment (HETDEX) Ly emitter candidates observed by DESI (landriau25a). |
| DESIVAST | Cosmic voids identified within the DESI DR1 volume (rincon25a). |
@@
| EmFit | Emission-line fitting results for galaxies (pucha25a). |
| FastSpecFit | Spectrophotometric fitting results from the FastSpecFit stellar continuum and emission-line modeling code. |
| Extended Halo-based Group | Halo-based group catalog based on Legacy Surveys DR9 for -band apparent magnitude galaxies (yang21a). |
| Mass EMLines | Stellar mass and emission line measurements for galaxies in DR1 (zou24b). |
@@
| AGN/Galaxy Classification | AGN and QSO identifications for galaxies from all target classes in DESI DR1 (Juneau et al. 2026, in preparation). |
| AGN Host Properties | Stellar masses and other physical properties from spectral energy distribution modeling which includes AGN templates (siudek24a). |
| BHMass | Iron-corrected supermassive black hole masses based on Mg II at (pan25a). |
| C IV Absorbers | Catalog of C IV absorber systems in DESI quasars (anand25a). |
| DLA NN and GP Finder | DLA parameters and detections using the NN and GP DLA finders (wang22a; ho20a). |
| DLA Template Finder | DLA parameters and detections using DLA Toolkit (brodzeller25a). |
| MgII Absorber | Summarized information of Mg II absorption systems in DESI quasars (napolitano23a). |
| ZLyA | Updated redshifts and BAL information used in the Ly Y1 BAO analysis (DESI Collaboration et al., 2025b). |
@@
| Ly Forest Y1 Deltas | Measured flux-transmission field used in the Ly Y1 BAO analysis (DESI Collaboration et al., 2025b). |
| Ly Forest Y1 Correlations | Measured correlations, distortion matrices and covariances used for the Ly Y1 BAO analysis (DESI Collaboration et al., 2025b). |
IV Data Access
In this section we provide a brief overview of how to access DR1. At the time of this writing, DESI catalogs and spectra can be accessed via file download (§IV.1) or a searchable database (§IV.2). We demonstrate these access methods, as well as techniques for exploring and manipulating DESI data, through a variety of tutorials and examples (§IV.3). The DESI collaboration may deploy additional data-access methods in the future, such as web services designed for interactive data exploration or interfaces enabling users to download individual spectra or custom collections of spectra. The latest information on how to access DESI data can always be found in the DESI documentation.282828https://data.desi.lbl.gov/doc/access